what is the perimeter of a right triangle whose shortest side is 3
Answer by CharlesG2(834)

what is the perimeter of a right triangle whose shortest side is 3
call other leg b
call hypotenuse c
c^2 = 3^2 + b^2 = 9 + b^2
P perimeter = 3 + b + sqrt(9 + b^2)